# Discovery Prospects for a Minimal Dark Matter Model at Cosmic and Intensity Frontier Experiments

### v1

@@ 3.2 Direct Detection
|¯$|\bar{\mathcal{M}}^{NR}(E_{R})|^{2}$ $=A^{2}F^{2}(E_{R})\left(\frac{f_{Z_{D}}}{m_{Z_{D}}^{2}+2m_{N}E_{R}}+\frac{f_{Z}}{m_{Z}^{2}}\right)^{2},$ (23)
@@
In the second line of Eq. 23, we defined a reference per-nucleon matrix element-squared |^$|\hat{\mathcal{M}}(E_{R})|^{2}$, as it will be useful in the following to separate the recoil energy dependence in the nuclear form factor from that in the partonic interaction.
@@ 3.2.1 Direct-Detection Constraints to Beam-Dump Parameter Space
Figure 6: The successive visualization of direct detection (DD) constraints, combined with indirect detection (ID) exclusions, in the dark photon parameter space. Top left: contours corresponding to the maximal allowed DD cross-section are drawn for fixed DM masses using the relationship between $\alpha_{D}$ and $\epsilon$ determined by the relic abundance curves, with colors corresponding to different regimes as indicated. Contours are labeled with DM mass in GeV. We require $m_{Z_{D}}<0.1m_{\chi}$. Regions above and to the left of each contour are excluded for the corresponding DM mass. Solid, dashed, and dotted lines indicate $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.9$, 0.7, and 0.5, respectively, as discussed in the text. Top right: we impose CMB constraints set by Eq. 18. Bottom left: we add the current and projected constraints from beam-dump experiments, as well as projections for LHCb. Additionally the orange line shows the neutrino fog lines for a 100 GeV DM particle. Bottom right: We compare the DD contours shown in the top left panel for DM masses in the range $3\,\mathrm{GeV}<m_{\chi}<12\,\mathrm{GeV}$ with several contours obtained using XENONnT results for DM of the same masses scattering through light mediators XENON:2024hup .
@@
The bound is shown in the top left panel of Fig. 6 for DM masses in the range $(0.5,100)$ GeV. The line weights of the curves are solid for regions with $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.9$, dashed for $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.7$, and dotted for all other regions. Since the DD cross section increases with increasing $\epsilon$ and decreasing $m_{Z_{D}}$, regions above and to the left of the curve are excluded for the corresponding DM mass. Here the blue portion of the curve indicates the freeze-in regime, while the green portion of the curve denotes the out-of-equilibrium freeze-out regime. These two different regimes predict different values of $\alpha_{D}(\epsilon)$, and thus must be independently constrained, i.e., the freeze-in regime is not excluded by the out-of-equilibrium freeze-out constraint. In the top right panel of Fig. 6 we additionally impose the CMB constraints derived in the previous subsection. These constraints eliminate a large fraction of the curves in the out-of-equilibrium freeze-out regime for lower dark matter masses. In the bottom left panel, we superimpose the constraints on the dark photon parameter space from beam dump and supernova bounds (gray regions) and indicate the projected sensitivity of future (beam dump and LHCb) experiments. We can conclude that visibly-decaying dark photons, consistent with current DD constraints, are discoverable at beam dump experiments. Note that in the regions relevant for future beam dump experiments, the direct detection exclusions in the out-of-equilibrium freeze-out regime are well characterized by the constant cross-section approximation. Also in this bottom left panel, we show a visualization of the neutrino fog in beam dump parameter space, indicated in orange. We use the neutrino fog cross-section for a 100 GeV DM particle scattering off of xenon, with the LZ recoil energy thresholds, taken from OHare:2021utq , and map it into dark photon parameter space following the same procedure as for the exclusions. We show only portions of this line that satisfy $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.7$, and in fact almost all of the out-of-equilibrium portion of the line satisfies $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.9$. Regions of parameter space below and to the right of the orange line lie below the neutrino fog, hence are challenging to access in direct detection experiments. We show the fog for a 100 GeV DM particle: this is the heaviest DM mass we consider, and it accordingly yields the maximal amount of parameter space above the neutrino fog in most of parameter space (although for lighter DM masses, regions of interest above the fog can extend to smaller values of $m_{Z_{D}}$ and $\epsilon$ in the out-of-equilibrium freeze-out regime). As evident in the figure, there is a notable chunk of parameter space in the out-of-equilibrium freeze-out regime that is inaccessible to intensity frontier experiments, but yields direct detection signals above the neutrino fog.
@@ 3.2.2 Beam-Dump Reach in Direct-Detection Parameter Space
We can visualize the reach of searches for visibly-decaying dark photons in the DD plane via a two-step procedure. First, we use the relic abundance constraint to determine $\alpha_{D}(\epsilon)$ for each value of $m_{\chi}$. We find it useful to separately analyze the freeze-in region $\alpha_{D}(\epsilon)<\alpha_{D}(\epsilon_{\mathrm{min}})$, the out-of-equilibrium freeze-out region with $\alpha_{D}^{TF}>\alpha_{D}(\epsilon)>\alpha_{D}(\epsilon_{\mathrm{min}})$, as well as the region above the thermalization floor, $\alpha_{D}(\epsilon)=\alpha_{D}^{TF}$.
@@
As can be seen from Fig. 3, the sensitivity of any given beam dump experiment and LHCb extends to some minimum value of $\epsilon$. This minimum value of epsilon is substantially above the value that separates the freeze-in and freeze-out regimes. Thus at a specific value of $m_{\chi}$, there is a gap between the values of $\alpha_{D}$ attained in the freeze-in region and those attained in the freeze-out regimes. This gap accounts for the separations of the freeze-in regions from the other regions in Fig. 7 (most notable for LHCb). The gap visible within each freeze-in region comes from the $\rho$-window cut. [^5] This gap is not visible in the out-of-equilibrium freeze-out region, since in this region the DD cross section and the DM relic abundance depend on different combinations of parameters.
@@
  • The combination of direct and indirect detection constraints rule out the above-the-thermalization floor scenario for all experiments. The panel for the DUNE experiment does not show any regions above the thermalization floor, since the DUNE reach is completely below the thermalization floor for the considered masses, as seen in Fig. 3.
@@ 4 Discussion and Conclusions
In this paper, we analyzed a secluded DM model where the DM is a Dirac fermion that interacts with the SM through the exchange of a dark photon. Within the considered area of parameter space, we find that the regime where the hidden sector is in thermal equilibrium with the SM at the time of DM freeze-out (the “WIMP-next-door” regime) is now entirely excluded, from a combination of CMB and direct detection constraints. However, there is still interesting parameter space for this model below the thermalization floor. This parameter space can be broadly split into two regimes: one where the thermal history of the model is dominated by an interplay of DM annihilations to dark photons with the out-of-equilibrium production of DM pairs from the SM, which we refer to here as “out-of-equilibrium freeze-out”, and one where the DM relic abundance is fixed through freeze-in from the SM.
@@
One important result of this work is that a discovery of dark photon that is the dark matter mediator is possible and consistent at future beam-dump experiments. In particular, this model still has viable parameter space for the proposed experiments DUNE, SHiP, and DarkQuest. In a large portion of this parameter space the DM attains its relic abundance through the relatively simple freeze-in process, but there are also surviving regions where the more involved out-of-equilibrium freeze-out scenario controls the DM abundance. A discovery could be made At LHCb, but only for dark matter in the freeze-in region.
@@
Meanwhile for nuclear-recoil direct detection experiments, the remaining discovery sensitivity to this specific model lies in the out-of-equilibrium freeze-out regime. In particular, there remains open parameter space in this regime with DD cross-sections above the neutrino fog that is inaccessible to planned beam-dump experiments. In this part of parameter space, future DD results offer the best discovery prospects for this model. Where the dark photons are heavy enough that their mass dominates over the recoil energy in DM-nucleon scattering, the bulk of the open model parameter space does lie below the neutrino fog, however, and in regions where the relic abundance is obtained through the simpler freeze-in process. Meanwhile for lighter dark photons, we have used XenonNT results for light mediators XENON:2024hup to show that, within the context of this minimal model, such light-mediator DD searches are generically sensitive to portions of dark photon parameter space that are already ruled out.
@@
Since potential connections to dark matter provide some of the best theoretical motivations for MeV-GeV scale dark photons, it is important to understand the potential cosmological consequences of discovering a visibly-decaying dark photon. Here we have shown that dark photon discovery at the next generation of intensity frontier experiments is consistent with a very minimal model of thermal dark matter. It is worth commenting that this minimal model admits a range of extensions that further expand its parameter space. For example, the model could contain a dark Higgs boson. In fact, in this paper, we invoked a Stückelberg origin for the dark photon mass, but the phenomenology of the model in the regime of interest does not depend strongly on this choice, and is also compatible with a dark Higgs mechanism. In particular, the DM in this model already has an $s$-wave annihilation cross-section and a leading spin- and velocity-independent contribution to the DM-nucleon cross-section. Thus the predictions of the minimal model are largely unchanged by the addition of a physical dark Higgs boson in the spectrum, provided that the vector portal coupling to the SM dominates the energy injection into the hidden sector compared to any Higgs-portal coupling resulting from dark Higgs-SM Higgs mixing. Cosmologically, requiring the vector portal to dominate over the Higgs portal can lead to problematically long dark Higgs boson lifetimes if the dark Higgs has only SM decay modes available. The resulting disruption to BBN can readily be avoided by ensuring that the dark Higgs is heavy enough to decay to pairs of dark photons.
@@
More consequentially, adding a small $U(1)_{D}$-breaking contribution to the DM mass splits the Dirac DM into two Majorana states, of which only the lightest is absolutely stable (see, e.g., Tucker-Smith:2001myb ; Alexander:2016aln ; CarrilloGonzalez:2021lxm ). This inelastic deformation can parametrically suppress both direct and indirect detection cross-sections, which will generally serve to enhance the relative discovery potential of visible dark photon searches.
@@ Acknowledgements.
We thank Julian Muñoz, Tracy Slatyer, and Jesse Thaler for useful conversations. The work of AA and JS was supported in part by U.S. Department of Energy (DOE) Office of High Energy Physics under grant numbers DE-SC0023365 and DE-SC0015655. The work of AA was supported in part by Kuwait University. The work of CC is supported in part by DOE grant DE-SC0012567 and in part by the Simons Foundation (Grant Number 929255, T.R.S). The research of SG is supported in part by DOE grant number DE-SC0010107. JS additionally thanks the Massachusetts Institute of Technology as well as the Kavli Institute for Theoretical Physics (KITP), supported in part by grant NSF PHY-2309135, for hospitality during the performance of this work.
@@ Appendix A Cross Section Expressions
$\mathcal{A}_{\chi\bar{\chi}\rightarrow Z_{D}Z_{D}}(s)=\frac{8\alpha_{D}}{\sqrt{s}}\sqrt{s-4m_{f}^{2}}\left(2\frac{(s^{2}+4sm_{\chi}^{2}-8m_{\chi}^{4})}{s(s-4m_{\chi}^{2})}\tanh^{-1}\left(\sqrt{\frac{s-4m_{\chi}^{2}}{s}}\right)\right).$ (32)

### v2

@@ 3.2 Direct Detection
$|\bar{\mathcal{M}}^{NR}(E_{R})|^{2}$ $=A^{2}F^{2}(E_{R})\left(\frac{f_{Z_{D}}}{m_{Z_{D}}^{2}+2m_{N}E_{R}}+\frac{f_{Z}}{m_{Z}^{2}}\right)^{2},$ (23)
@@
In the second line of Eq. 23, we defined a reference per-nucleon matrix element-squared $|\hat{\mathcal{M}}(E_{R})|^{2}$, as it will be useful in the following to separate the recoil energy dependence in the nuclear form factor from that in the partonic interaction.
@@ 3.2.1 Direct-Detection Constraints to Beam-Dump Parameter Space
Figure 6: The successive visualization of direct detection (DD) constraints, combined with indirect detection (ID) exclusions, in the dark photon parameter space. Top left: contours corresponding to the maximal allowed DD cross-section are drawn for fixed DM masses using the relationship between $\alpha_{D}$ and $\epsilon$ determined by the relic abundance curves, with colors corresponding to different regimes as indicated. Contours are labeled with DM mass in GeV. We require $m_{Z_{D}}<0.1m_{\chi}$. Regions above and to the left of each contour are excluded for the corresponding DM mass. Thick, medium, and thin line weights indicate $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.9$, 0.7, and 0.5, respectively, as discussed in the text. Top right: we impose CMB constraints set by Eq. 18. Bottom left: we add the current and projected constraints from SN1987a and beam-dump experiments, as well as projections for LHCb. Additionally the orange line shows the neutrino fog lines for a 100 GeV DM particle. Bottom right: We compare the DD contours shown in the top left panel for DM masses in the range $3\,\mathrm{GeV}<m_{\chi}<12\,\mathrm{GeV}$ with several contours obtained using XENONnT results for DM of the same masses scattering through light mediators XENON:2024hup .
@@
The bound is shown in the top left panel of Fig. 6 for DM masses in the range $(0.5,100)$ GeV. The line weights of the curves are thick for regions with $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.9$, medium for $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.7$, and thin for all other regions. Since the DD cross section increases with increasing $\epsilon$ and decreasing $m_{Z_{D}}$, regions above and to the left of the curve are excluded for the corresponding DM mass. Here the blue portion of the curve indicates the freeze-in regime, while the green portion of the curve denotes the out-of-equilibrium freeze-out regime. These two different regimes predict different values of $\alpha_{D}(\epsilon)$, and thus must be independently constrained, i.e., the freeze-in regime is not excluded by the out-of-equilibrium freeze-out constraint. In the top right panel of Fig. 6 we additionally impose the CMB constraints derived in the previous subsection. These constraints eliminate a large fraction of the curves in the out-of-equilibrium freeze-out regime for lower dark matter masses. In the bottom left panel, we superimpose the constraints on the dark photon parameter space from beam dump and supernova bounds (gray regions) and indicate the projected sensitivity of future (beam dump and LHCb) experiments. We can conclude that visibly-decaying dark photons, consistent with current DD constraints, are discoverable at beam dump experiments. Note that in the regions relevant for future beam dump experiments, the direct detection exclusions in the out-of-equilibrium freeze-out regime are well characterized by the constant cross-section approximation. Also in this bottom left panel, we show a visualization of the neutrino fog in beam dump parameter space, indicated in orange. We use the neutrino fog cross-section for a 100 GeV DM particle scattering off of xenon, with the LZ recoil energy thresholds, taken from OHare:2021utq , and map it into dark photon parameter space following the same procedure as for the exclusions. We show only portions of this line that satisfy $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.7$, and in fact almost all of the out-of-equilibrium portion of the line satisfies $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}>0.9$. Regions of parameter space below and to the right of the orange line lie below the neutrino fog, hence are challenging to access in direct detection experiments. We show the fog for a 100 GeV DM particle: this is the heaviest DM mass we consider, and it accordingly yields the maximal amount of parameter space above the neutrino fog in most of parameter space (although for lighter DM masses, regions of interest above the fog can extend to smaller values of $m_{Z_{D}}$ and $\epsilon$ in the out-of-equilibrium freeze-out regime). As evident in the figure, there is a notable chunk of parameter space in the out-of-equilibrium freeze-out regime that is inaccessible to intensity frontier experiments, but yields direct detection signals above the neutrino fog.
@@ 3.2.2 Beam-Dump Reach in Direct-Detection Parameter Space
We can visualize the reach of searches for visibly-decaying dark photons in the DD plane via a two-step procedure. First, we use the relic abundance constraint to determine $\alpha_{D}(\epsilon)$ for each value of $m_{\chi}$. We find it useful to separately analyze the freeze-in region $\alpha_{D}(\epsilon)<\alpha_{D}(\epsilon_{\mathrm{min}})$, the out-of-equilibrium freeze-out region with $\alpha_{D}(\epsilon_{\mathrm{min}})<\alpha_{D}(\epsilon)<\alpha_{D}^{TF}$, as well as the region above the thermalization floor, $\alpha_{D}(\epsilon)=\alpha_{D}^{TF}$.
@@
As can be seen from Fig. 3, the sensitivity of any given beam dump experiment and LHCb extends to some minimum value of $\epsilon$. This minimum value of $\epsilon$ is substantially above the value that separates the freeze-in and freeze-out regimes. Thus at a specific value of $m_{\chi}$, there is a gap between the values of $\alpha_{D}$ attained in the freeze-in region and those attained in the freeze-out regimes. This gap accounts for the separations of the freeze-in regions from the other regions in Fig. 7 (most notable for LHCb). The gap visible within each freeze-in region comes from the $\rho$-window cut. [^6] This gap is not visible in the out-of-equilibrium freeze-out region, since in this region the DD cross section and the DM relic abundance depend on different combinations of parameters.
@@
  • The combination of direct and indirect detection constraints rule out the above-the-thermalization floor scenario for all experiments. As already seen in Fig. 3, the parameter space probed by DUNE, DarkQuest, and SHiP lies below the thermalization floor for the considered masses. Here we also see that in the parameter space probed by LHCb, direct detection constraints leave freeze-in as the only viable regime.
@@ 4 Discussion and Conclusions
In this paper, we analyzed a secluded DM model where the DM is a Dirac fermion that interacts with the SM through the exchange of a dark photon. We focus on the DM mass range GeV $\lesssim m_{\chi}\lesssim 100$ GeV: light enough that DM production in the early universe is dominated by temperatures below electroweak symmetry breaking, and heavy enough for nuclear recoil, rather than electron recoil DAMIC-M:2025luv , to dominate direct detection constraints. Within the considered area of parameter space, we find that regions above the thermalization floor (the “WIMP-next-door” regime) are now entirely excluded through a combination of CMB and direct detection constraints. However, there is still interesting parameter space for this model below the thermalization floor. This parameter space can be broadly split into two regimes: one where the thermal history of the model is dominated by an interplay of DM annihilations to dark photons with the out-of-equilibrium production of DM pairs from the SM, which we refer to here as “out-of-equilibrium freeze-out”, and one where the DM relic abundance is fixed through freeze-in from the SM.
@@
One important result of this work is that a discovery of a visibly-decaying dark photon at future beam-dump experiments is consistent with this very minimal model of dark matter. In particular, this model still has viable parameter space for the proposed experiments DUNE, SHiP, and DarkQuest. In a large portion of this parameter space the DM attains its relic abundance through the relatively simple freeze-in process, but there are also surviving regions where the more involved out-of-equilibrium freeze-out scenario controls the DM abundance. A discovery could be made at LHCb, but only for dark matter in the freeze-in region. If a dark photon were discovered at a collider experiment, the model would also predict a small but non-zero dark matter pair production cross section, which given the dark photon parameters would be determined up to the dark matter mass. This signature would be a challenging but definite target for the LHC and future colliders, which would provide complementary information on the connection between the observed dark photon and dark matter. Similarly, a dark photon discovery would provide motivation to pursue DD sensitivity below the neutrino fog.
@@
Meanwhile for nuclear-recoil direct detection experiments, the remaining discovery sensitivity to this specific model lies primarily in the out-of-equilibrium freeze-out regime. In particular, there remains open parameter space in this regime, with DD cross-sections above the neutrino fog, that is inaccessible to planned beam-dump experiments. In this part of parameter space, future DD results offer the best discovery prospects for this model. Where the dark photons are heavy enough that their mass dominates over the recoil energy in DM-nucleon scattering, the bulk of the open model parameter space does lie below the neutrino fog, however. Meanwhile for lighter dark photons, we have used XenonNT results for light mediators XENON:2024hup to show that, within the context of this minimal model, such light-mediator DD searches are generically sensitive to portions of dark photon parameter space that are already ruled out.
It is worth bearing in mind that all the terrestrial and astrophysical constraints on visibly-decaying dark photons that we have quoted are subject to uncertainties. At beam dumps, updated calculations of and systematic uncertainties in various dark photon production processes can notably impact experimental sensitivities Curtin:2023bcf ; Kyselov:2024dmi ; Zhou:2024aeu . Supernova constraints, meanwhile, are sensitive to uncertainties in the detailed modeling of the progenitor, in particular its mass. Additionally, constraints on energy deposition in the stellar envelope from BSM particles can strengthen the supernova constraints on the production of light feebly-interacting particles Falk:1978kf ; Sung:2019xie ; Shin:2022ulh ; Caputo:2022mah . The resulting constraints from SN1987a computed in Sung:2019xie for dark photons are accordingly somewhat more stringent than what we show here given their fiducial modeling of the progenitor mass. In particular these bounds extend the excluded area of the dark photon parameter space in the region where the dark photon is appreciably light in direct detection experiments, i.e., for $\sigma_{\chi n}^{\mathrm{eff}}/\sigma^{0}_{\chi n}\lesssim 0.5$. However our main conclusions about the complimentary discovery sensitivity of both beam dump and DD experiments are not affected by the uncertainties in current constraints on visibly-decaying dark photons.
@@
Since potential connections to dark matter provide some of the best theoretical motivations for MeV-GeV scale dark photons, it is important to understand the potential cosmological consequences of discovering a visibly-decaying dark photon. Here we have shown that dark photon discovery at the next generation of intensity frontier experiments is consistent with a very minimal model of thermal dark matter. It is worth commenting that this minimal model admits a range of extensions that further expand its parameter space. For example, the model could contain a dark Higgs boson. In this paper, we invoked a Stückelberg origin for the dark photon mass, but the phenomenology of the model in the regime of interest does not in fact depend strongly on this choice, and is also compatible with a dark Higgs mechanism. In particular, the DM in this model already has an $s$-wave annihilation cross-section and a leading spin- and velocity-independent contribution to the DM-nucleon cross-section. Thus the predictions of the minimal model are largely unchanged by the addition of a physical dark Higgs boson in the spectrum, provided that the vector portal coupling to the SM dominates the energy injection into the hidden sector compared to any Higgs-portal coupling resulting from dark Higgs-SM Higgs mixing. Cosmologically, requiring the vector portal to dominate over the Higgs portal can lead to problematically long dark Higgs boson lifetimes if the dark Higgs has only SM decay modes available. The resulting disruption to BBN can readily be avoided by ensuring that the dark Higgs is heavy enough to decay to pairs of dark photons.
@@
More consequentially, adding a small $U(1)_{D}$-breaking contribution to the DM mass splits the Dirac DM into two Majorana states, of which only the lightest is absolutely stable (see, e.g., Tucker-Smith:2001myb ; Alexander:2016aln ; CarrilloGonzalez:2021lxm ). This inelastic deformation can parametrically suppress direct detection cross-sections, which will generally serve to enhance the relative discovery potential of visible dark photon searches, while introducing additional cosmological signatures in the freeze-in regime Heeba:2023bik .
@@ Acknowledgements.
We thank Julian Muñoz, Tracy Slatyer, Katelin Schutz, Jesse Thaler, and Douglas Tuckler for useful conversations. The work of AA and JS was supported in part by U.S. Department of Energy (DOE) Office of High Energy Physics under grant numbers DE-SC0023365 and DE-SC0015655. The work of AA was supported in part by Kuwait University. The work of CC is supported in part by DOE grant DE-SC0012567 and in part by the Simons Foundation (Grant Number 929255, T.R.S). The research of SG is supported in part by DOE grant number DE-SC0010107. JS additionally thanks the Massachusetts Institute of Technology as well as the Kavli Institute for Theoretical Physics (KITP), supported in part by grant NSF PHY-2309135, for hospitality during the performance of this work.
@@ Appendix A Cross Section Expressions
$\mathcal{A}_{\chi\bar{\chi}\rightarrow Z_{D}Z_{D}}(s)=8\pi\alpha_{D}^{2}\sqrt{1-\frac{4m_{\chi}^{2}}{s}}\left(2\frac{(s^{2}+4sm_{\chi}^{2}-8m_{\chi}^{4})}{s(s-4m_{\chi}^{2})}\tanh^{-1}\left(\sqrt{\frac{s-4m_{\chi}^{2}}{s}}\right)-\frac{(s+4m_{\chi}^{2})}{\sqrt{s(s-4m_{\chi}^{2})}}\right).$ (32)